For the network of capacitors shown, find the equivalent capacitance and the total charge when connected to a 36 V source. Note that C1 = C2 = 4.0 F and C3 = C4 = 6.0 F.



1
Expert's answer
2022-04-11T16:55:32-0400

"C_{12}=C_1+C_2=8~F,"

"C_{34}=C_3+C_4=12~F,"

"C=\frac{C_{12}C_{34}}{C_{12}+C_{34}}=4.8~F,"

"q=CU=173~C."